# VoiceBench: Benchmarking LLM-Based Voice Assistants

### v1

@@ Abstract
Building on the success of large language models (LLMs), recent advancements such as GPT-4o have enabled real-time speech interactions through LLM-based voice assistants, offering a significantly improved user experience compared to traditional text-based interactions. However, the absence of benchmarks designed to evaluate these speech interaction capabilities has hindered progress of LLM-based voice assistants development. Current evaluations focus primarily on automatic speech recognition (ASR) or general knowledge evaluation with clean speeches, neglecting the more intricate, real-world scenarios that involve diverse speaker characteristics, environmental and content factors. To address this, we introduce VoiceBench, the first benchmark designed to provide a multi-faceted evaluation of LLM-based voice assistants. VoiceBench also includes both real and synthetic spoken instructions that incorporate the above three key real-world variations. Extensive experiments reveal the limitations of current LLM-based voice assistant models and offer valuable insights for future research and development in this field. [^1] Data is available at https://github.com/MatthewCYM/VoiceBench.
@@ 3.1 Dataset Construction
To evaluate the general knowledge of voice assistants, we incorporate three datasets: AlpacaEval, CommonEval, and SD-QA. For AlpacaEval (Li et al., 2023), we follow LLaMA-Omni (Fang et al., 2024) by retaining the helpful_base and vicuna subsets, and removing questions related to mathematics and coding. For CommonEval, we manually collect information-seeking questions from CommonVoice (Ardila et al., 2020), which features speech recorded in realistic settings by a diverse range of speakers using their personal devices. SD-QA (Faisal et al., 2021) consists of spoken questions with varying accents from the TyDi-QA dataset (Clark et al., 2020). In its original format, models are required to answer these questions using context passages. However, in voice interaction scenarios, providing extensive context is often impractical. Therefore, we feed only the spoken questions to voice assistants, requiring them to respond using their internal knowledge. Upon reviewing the dataset, we observe that some questions, such as "How many people live in Dallas?", require up-to-date information that frequently changes and cannot be answered without additional context. To ensure fair evaluation, we retain only those questions that could be answered using the internal knowledge of large language models (LLMs). Specifically, we prompt three advanced LLMs—Claude-3.5 (Anthropic, 2024)222claude-3-5-sonnet@20240620, Gemini-1.5-Pro (Reid et al., 2024)333gemini-1.5-pro-002, and GPT-4o (Achiam et al., 2023)444gpt-4o-2024-08-06—to answer these questions, and we manually select those questions that received correct answers from at least one of the models.
@@
The data statistics of proposed VoiceBench is summarized in Tab. 1. CommonEval and SD-QA contain real spoken instructions, while AlpacaEval, IFEval, and AdvBench feature synthetic spoken instructions. All instructions are relatively short, making them well-suited for voice interactions. These datasets are designed to evaluate the multi-faceted capabilities of voice assistants.
@@
Table 1: Data statistics of VoiceBench.
@@ 3.2 Experiment Setup
We evaluate various end-to-end voice assistants on the proposed VoiceBench, including Qwen2-Audio (Chu et al., 2024), LLaMA-Omni (Fang et al., 2024), Mini-Omni (Xie and Wu, 2024a), VITA (Fu et al., 2024), and DiVA (Held et al., 2024). Additionally, we build a naive voice assistant pipeline, where an automatic speech recognizer transcribes the input speech query into text, and a text-only LLM generates a response based on the transcribed query. The architectures of the evaluated models are summarized in Tab. 2.
@@
Naïve 4.81 4.68 4.38 4.04 77.76 75.41 72.33 68.54 73.91 79.52 65.37 73.70 96.54 98.08 80.27 74.50
DiVA 4.84 3.86 4.29 3.54 78.30 74.50 62.39 51.72 68.70 76.31 34.93 43.38 99.23 98.27 81.14 64.02
LLaMA-Omni 4.58 3.95 4.32 3.46 55.33 60.40 40.14 39.24 45.38 56.53 10.15 19.58 98.46 11.35 69.26 40.21
Mini-Omni 2.64 2.25 2.55 2.02 26.04 7.23 23.69 4.16 13.04 22.89 8.99 18.17 86.35 37.12 57.11 41.56
Qwen2-Audio 4.27 3.89 3.77 3.43 61.66 40.69 41.77 29.66 28.70 38.06 20.73 31.93 96.73 96.73 72.72 59.83
VITA 4.16 3.78 3.88 2.15 72.69 76.13 31.28 24.59 48.99 57.53 18.12 27.51 95.19 26.73 75.14 39.33
Table 3: The performance of various voice assistants on VoiceBench. The Text and Speech columns refer to the model performance with text-form and speech-form instructions respectively. We report the performance on SD-QA United States accent above. The best performance on each task is bold.
@@ 3.3 Results
Firstly, the naive pipeline-based voice assistant significantly outperforms all end-to-end models on spoken instructions, with a margin exceeding 10 points. Among the end-to-end models, Mini-Omni underperforms significantly due to its use of a smaller speech encoder and base LLM, which prioritize processing efficiency over performance. Notably, while some end-to-end models, such as Qwen2-Audio (Chu et al., 2024), demonstrate comparable or even superior ASR and audio analysis performance Yang et al. (2024) compared to models such as Whisper (Radford et al., 2023) or the naive pipeline, a significant performance gap persists when handling spoken instructions in our AudioBench evaluation. This highlights a major limitation in current voice assistant evaluations, which rely heavily on ASR metrics.
@@
Thirdly, we observe a notable disparity between the text and speech processing abilities of current models. The naive pipeline model shows a relatively small performance gap of 5.77 points from text to speech instructions, primarily due to ASR errors introduced by the speech recognition sub-model within the pipeline. In contrast, end-to-end models exhibit much larger gaps. For instance, VITA demonstrates a performance gap exceeding 35 points when taking text and speech inputs.

### v2

@@ Abstract
With the advent of large language models (LLMs), recent advancements such as GPT-4o have enabled real-time speech interactions through LLM-based voice assistants, offering a significantly improved user experience over traditional text-based interactions. This calls for an appropriate benchmark designed to evaluate such speech interactions. The prior studies on related evaluations focus primarily on automatic speech recognition (ASR) or general knowledge evaluation with clean speeches. There exists a disconnect with the intricate, real-world scenarios that involve diverse speaker characteristics, environmental and content factors. To address this, we introduce VoiceBench, the first benchmark designed to provide a multi-faceted evaluation of LLM-based voice assistants. VoiceBench also includes both real and synthetic spoken instructions that incorporate the above three key real-world variations. Extensive experiments reveal the limitations of current LLM-based voice assistant models and offer valuable insights for future research and development in this field. [^1] Code and data is available at https://github.com/MatthewCYM/VoiceBench.
@@ 3.1 Dataset Construction
To evaluate the general knowledge of voice assistants, we incorporate three datasets: AlpacaEval, CommonEval, and SD-QA. For AlpacaEval (Li et al., 2023), we follow LLaMA-Omni (Fang et al., 2024) to remove questions related to mathematics and coding. For CommonEval, we manually collect information-seeking questions from CommonVoice (Ardila et al., 2020), which features speech recorded in realistic settings by a diverse range of speakers using their personal devices. SD-QA (Faisal et al., 2021) consists of spoken questions with varying accents from the TyDi-QA dataset (Clark et al., 2020). In its original format, models are required to answer these questions using context passages. However, in voice interaction scenarios, providing extensive context is often impractical. Therefore, we feed only the spoken questions to voice assistants, requiring them to respond using their internal knowledge. Upon reviewing the dataset, we observe that some questions, such as "How many people live in Dallas?", require up-to-date information that frequently changes and cannot be answered without additional context. To ensure fair evaluation, we retain only those questions that could be answered using the internal knowledge of large language models (LLMs). Specifically, we prompt three advanced LLMs—Claude-3.5 (Anthropic, 2024)222claude-3-5-sonnet@20240620, Gemini-1.5-Pro (Reid et al., 2024)333gemini-1.5-pro-002, and GPT-4o (Achiam et al., 2023)444gpt-4o-2024-08-06—to answer these questions, and we manually select those questions that received correct answers from at least one of the models.
@@
The data statistics of proposed VoiceBench is summarized in Tab. 1. CommonEval and SD-QA contain real spoken instructions, while AlpacaEval, IFEval, and AdvBench feature synthetic spoken instructions. All instructions are relatively short, making them well-suited for voice interactions. These datasets are designed to evaluate the multi-faceted capabilities of voice assistants. Additionally, we present the evaluation results on the full AlpacaEval in Sec. 3.3. For all other experiments, we follow LLaMA-Omni (Fang et al., 2024) to use the helpful_base and vicuna subsets only.
@@
Table 1: Data statistics of VoiceBench. AlpacaEval includes helpful_base and vicuna subsets.
@@ 3.2 Experiment Setup
We evaluate various end-to-end voice assistants on the proposed VoiceBench, including Qwen2-Audio (Chu et al., 2024), LLaMA-Omni (Fang et al., 2024), Mini-Omni (Xie and Wu, 2024a), Mini-Omni2 (Xie and Wu, 2024b), VITA (Fu et al., 2024), and DiVA (Held et al., 2024). Additionally, we build a naive voice assistant pipeline, where an automatic speech recognizer transcribes the input speech query into text, and a text-only LLM generates a response based on the transcribed query. The architectures of the evaluated models are summarized in Tab. 2.
@@
Naive 4.69 4.53 4.38 4.04 77.76 75.41 72.33 68.54 73.91 79.52 65.37 73.70 96.54 98.08 86.25 81.88
DiVA 4.68 3.67 4.29 3.54 78.30 74.50 62.39 51.72 68.70 76.31 34.93 43.38 99.23 98.27 85.51 67.73
LLaMA-Omni 4.39 3.70 4.32 3.46 55.33 60.40 40.14 39.24 45.38 56.53 10.15 19.58 98.46 11.35 76.29 41.83
Mini-Omni 2.34 1.95 2.55 2.02 26.04 7.23 23.69 4.16 13.04 22.89 8.99 18.17 86.35 37.12 43.75 28.80
Mini-Omni2 2.65 2.32 2.86 2.18 13.02 9.76 11.03 7.59 10.15 17.87 7.25 15.86 92.88 57.50 45.70 33.67
Qwen2-Audio 4.11 3.74 3.77 3.43 61.66 40.69 41.77 29.66 28.70 38.06 20.73 31.93 96.73 96.73 67.78 60.45
VITA 4.00 3.38 3.88 2.15 72.69 76.13 31.28 24.59 48.99 57.53 18.12 27.51 95.19 26.73 76.08 37.62
Table 3: The performance of various voice assistants on VoiceBench. The Text and Speech columns refer to the model performance with text-form and speech-form instructions respectively. We report the performance on SD-QA United States accent above.
@@ 3.3 Results
Firstly, the naive pipeline-based voice assistant significantly outperforms all end-to-end models on spoken instructions, with a margin exceeding 10 points. Among the end-to-end models, Mini-Omni and Mini-Omni2 underperform significantly due to its use of a smaller speech encoder and base LLM, which prioritize processing efficiency over performance. Notably, while some end-to-end models, such as Qwen2-Audio (Chu et al., 2024), demonstrate comparable or even superior ASR and audio analysis performance Yang et al. (2024) compared to models such as Whisper (Radford et al., 2023) or the naive pipeline, a significant performance gap persists when handling spoken instructions in our AudioBench evaluation. This highlights a major limitation in current voice assistant evaluations, which rely heavily on ASR metrics.
@@
Thirdly, we observe a notable disparity between the text and speech processing abilities of current models. The naive pipeline model shows a relatively small performance gap of 4.37 points from text to speech instructions, primarily due to ASR errors introduced by the speech recognition sub-model within the pipeline. In contrast, end-to-end models exhibit much larger gaps. For instance, VITA demonstrates a performance gap exceeding 35 points when taking text and speech inputs.